# Compelling ReLU Networks to Exhibit Exponentially Many Linear Regions at Initialization and During Training

### v4

@@ 1 Introduction
Beyond complementary advances in areas like hardware, storage, and networking, the success of neural networks is primarily due to their ability to efficiently capture and represent nonlinear functions (Gibou et al., 2019). In a neural network, the goal of an activation function is to introduce nonlinearity between the network’s layers so that the network does not simplify to a single linear function. The rectified linear unit (ReLU) has a unique interpretation in this regard. Since it either deactivates a neuron or acts as an identity, the resulting transformation on each individual input remains linear. However, each possible configuration of active and inactive neurons can produce a unique linear transformation over a particular region of input space. The number of these activation patterns and their corresponding linear regions provides a way to measure the expressivity of a ReLU network [^1] The reader is referred to Chmielewski-Anders (2020) for definitions of linear regions, activation patterns, and activation regions. and can theoretically scale exponentially with the depth of the network (Montufar et al., 2014a; Serra et al., 2018). Hence deep architectures may outperform shallow ones.
@@
The present work aims to begin eliminating these inefficiencies, starting in a simple one-dimensional setting. Drawing inspiration from existing theoretical ReLU constructions, our novel contributions include a special reparameterization of a ReLU network that forces it to maintain an exponential number of activation patterns over the input domain. We then demonstrate a novel pretraining strategy, which trains these derived parameters before manipulating the underlying matrix weights. This allows the network to discover solutions that are more accurate and unlikely to be found otherwise. Our technique directly rectifies the issues raised by Hanin & Rolnick (2019). Firstly, we initialize immediately with an exponential number of linear regions, which would not be expected to happen otherwise. Secondly, we are not reliant on gradient descent to “discover” new regions; we only need to maintain the existing ones, which is already guaranteed by the reparameterization during the pretraining stage. Additionally, this pretraining step can avoid areas of the loss landscape where unassisted gradient descent might make short-sighted optimizations that eliminate activation regions. Our results demonstrate that minimizing the reliance of network training on unassisted gradient descent can reliably produce error values orders of magnitude lower than a traditionally-trained network of equal size. Although the preliminary results in this theoretical study pertain to relatively low-dimensional functions, the paper concludes with our views on extending these theoretical exponential benefits for ReLU networks to arbitrary smooth functions with arbitrary dimensionality, which would have significant practical utility.
@@ 2 Related Work
This work is primarily concerned with a novel training methodology, but it also possesses a significant approximation theoretic component. In one dimension, where we initially focus, our work can be viewed as a first step towards generalizing the approximation to $x^{2}$ we review in this section. The reparameterization we employ modifies this method to become trainable in order to represent other convex one-dimensional functions, and then converts that result back into a matrix representation.
@@
2.1 Related Work in Approximation
Infinitely wide neural networks are known to be universal function approximators, even with only one hidden layer (Hornik et al., 1989; Cybenko, 1989). Infinitely deep networks of fixed width are universal approximators as well (Lu et al., 2017; Hanin, 2019). In finite cases, the trade-off between width and depth is often studied in assessing a network’s ability to approximate (learn) a function.
@@ 2.1 Related Work in Approximation
The dilated triangular waveforms produced in this manner are not particularly useful on their own. Their oscillations quickly become excessively rapid, and their derviatives do not exist everywhere (especially in the infinite depth limit). But these problems can be rectified by taking a sum over the layers of a network. Yarotsky (2017) and Liang & Srikant (2016) demonstrate the ability to construct $y=x^{2}$ on $[0,1]$ with exponential accuracy using symmetric triangle waves. To produce their approximation, one begins with $f_{0}(x)=x$, then computes $f_{1}(x)=f_{0}(x)-T(x)/4$, $f_{2}(x)=f_{1}(x)-T(T(x))/16$, $f_{3}(x)=f_{2}(x)-T(T(T(x)))/64$, and so forth, as pictured in Figure 1. As these successive approximations are computed, Figure 1 plots their convergence to $x^{2}$, as well as the convergence of the derivative to $2x$. Our reparameterization generalizes this approximation to use non-symmetric triangle waves to approximate a wider class of convex, differentiable, one-dimensional functions.
@@
The $x^{2}$ approximation is used by other theoretical works as a building block to guarantee exponential convergence rates in more complex systems. One possible use case is to construct a multiplication gate. Perekrestenko et al. (2018) does so via the identity $(x+y)^{2}=x^{2}+y^{2}+2xy$. The squared terms can all be moved to one side, expressing the product $xy$ as a linear combination of squared terms. They then further assemble these multiplication gates into an interpolating polynomial, which can have an exponentially decreasing error when the interpolation points are chosen to be the Chebyshev nodes. Polynomial interpolation does not scale well into high dimensions, so this and papers with similar approaches will usually come with restrictions that limit function complexity: Wang et al. (2018) requires low input dimension, Montanelli et al. (2020) uses band limiting, and Chen et al. (2019) approximates low-dimensional manifolds. These works all make use of a fixed representation of $x^{2}$. If our networks were substituted in for the $x^{2}$ approximation, these works would provide theoretical guarantees about the capabilities of the resulting model. Even though the approximation rates will not scale well with input dimension, they serve as a bound that can be improved upon. In Section 5 we further elaborate on how to use our networks to represent higher-dimensional or non-convex functions.
@@
Lastly, other works focus on showing how ReLU networks can encode and subsequently surpass traditional approximation methods (Lu et al., 2021; Daubechies et al., 2022). Interestingly, certain fundamental themes from above like composition, triangles, or squaring are still present. Another interesting comparison of the present work is to Ivanova & Kubat (1995), which uses decision trees as a means to initialize neural networks. It is a sigmoid/classification analogy to this work, but rather than an attempting to improve neural networks with decision trees, it is an attempt to improve decision trees with neural networks. Finally, Park et al. (2021) approaches the problem of linear region maximization from an information theory perspective and uses a loss penalty rather than a reparameterization to increase the number of linear regions.
2.2 Related Work in Initialization
@@
Our work seeks to improve network initialization by making use of explicit theoretical constructs. This stands in sharp contrast the current standard approach, which treats neurons homogeneously. Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot & Bengio, 2010). They use weight values that are sampled from distributions defined by the input and output dimension of each layer. Aside from sub-optimal approximation power associated with random weights, a common issue is that an entire ReLU network can collapse into outputting a constant value at initialization. This is referred to as the dying ReLU phenomenon (Qi et al., 2024; Nag et al., 2023). It occurs when the initial weights and biases cause every neuron in a particular layer to output a negative value. The ReLU activation then sets the output of that layer to 0, blocking any gradient updates. Worryingly, as depth goes to infinity, the dying ReLU phenomenon becomes increasingly likely (Lu et al., 2020). Several papers propose solutions: Shin & Karniadakis (2020) use a data-dependent initialization, while Singh & Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed. We observed during our experiments that RAAI greatly reduces, but does not eliminate the likelihood of dying ReLU. Our approach enforces a specific network structure that does not collapse in this manner.
@@ 3 Initialization and Pretraining Construction
Naively using the output of one triangle generator (from Figure 2) as input to the next would form a $1\times 2\times 1\times 2\times 1$ shape, but this is unnecessarily deep. We can replicate the one-dimensional function composition in the hidden layers on the left side of Figure 2 by using weight sharing instead. Any outgoing weight from $t_{1}$ or $t_{2}$ is shared; every neuron taking in a triangle wave as input does so by combining $t_{1}$ and $t_{2}$ in the same proportion. In this way, we can avoid having to use the extra intermediate neurons. This is evident in Figure 3 where the waves output by $t_{1}$ and $t_{2}$ neurons posses identical slopes in regions where both neurons are active.
@@ 3.1 Differentiable Model Output
Given the rapid oscillations of the triangle waves formed at each layer, the resulting output of the network will appear jagged or fractal with many choices of scaling parameters. This can lead to the network settling into solutions that are not predictive of the areas in between training points (or possibly not even interpolating the training data) due to rapid oscillation. Our main mathematical result addresses this issue by forcing differentiability of the network output. To achieve differentiability, it turns out that for this class of functions, the peak locations of the triangle waves uniquely determine the scales with which to sum them:
@@
$F^{\prime}(x)$ is defined on $[0,1]$ only if the scaling coefficients $s_{i}$ are selected based on the triangle peaks $a_{i}$ according to:
@@ 4 One-Dimensional Experiments
The goal of deep learning is to train a network to approximate a (generally unknown) nonlinear function. Accordingly, in this section, we implement our method to learn several convex one-dimensional curves. These nonlinear functions are known—so that we can measure error—and we remark on more difficult functions in Sections 5 and 6. The aim of these experiments is twofold: (1) we would like to determine how to learn the most effective function representations possible, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data. To demonstrate that our networks can learn function representations that better utilize depth, we benchmark against PyTorch’s (Paszke et al., 2019) default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh & Sreejith (2021), and produce errors that are orders of magnitude lower than both. For our experimental networks, we use a common set of initialization points where the triangle peaks and scaling parameters are chosen according to our main theorem (Equation 4). We compare the effect of reparameterized pretraining against skipping immediately to standard gradient descent (from a reparameterized initialization as in Section 3). We also compare training the scaling parameters freely during pretraining, instead of choosing them to achieve differentiability. The intent of both comparisons is see if pretraining and further differentiability constraints facilitate smoother navigation of the loss landscape. Lastly, we conduct a second round of tests to determine if pretrained networks display an enhanced predictive capacity on unseen data points, as might be expected if they can leverage greater nonlinearity in their outputs.
@@ 4.2 Numerical Results
Default Network $2.11\times 10^{-5}$ $2.19\times 10^{-5}$ $7.20\times 10^{-2}$ $2.82\times 10^{-2}$
@@
Default Network $4.50\times 10^{-5}$ $5.75\times 10^{-5}$ $1.15\times 10^{-1}$ $1.96\times 10^{-1}$
@@
Default Network $2.41\times 10^{-4}$ $2.14\times 10^{-3}$ $2.27\times 10^{-5}$ $1.60\times 10^{-4}$
@@
Our first set of results are shown in Tables 1 and 2, wherein we observe several important trends. First, the worst performing networks are those that rely on purely randomized initializations. Even the networks that forgo pretraining benefit from initializing with many activation regions. When pretraining constraints are used, they are able to steer gradient descent to the best solutions, resulting in reductions in minimum error of three orders of magnitude over default networks.
Pretraining with differentiability enforced also closes the gap between the minimum and mean errors compared to other setups. This indicates that these loss landscapes are indeed the most reliable to traverse. Enforcing differentiability during pretraining can impart a bias towards smoother solutions during subsequent unassisted gradient descent.
@@
Figure 4: Comparison between standard initialization/gradient descent and pretraining with differentiablity enforced. Using more linear regions allows the curve to better predict the test points.
@@ 5 Extension to Non-Convex Functions and Higher Dimensions
The same approach can address non-convex functions since many non-convex functions (including every continuous function with a bounded second derivative (Hessian)) can be expressed as the difference of two convex functions (see Zhang et al. (2018), which also describes how every ReLU network can be decomposed into a difference of two piecewise linear convex functions). We do not claim that two copies of our network are sufficient for learning every pathological function, but preliminary results in this section are surprisingly promising (and of course, assembling more copies of our network into a deeper or wider larger network would give additional representational power).
@@
Figures 7 and 8 show learning the non-convex function $y=x^{3}-x$ and the two-dimensional function $z=r^{3}$ on $[-1,1]$, each by using two of our convex function blocks. Rather than directly using our networks as activation functions, we instead encode each of their layers as a block diagonal structure in one combined matrix (which has the same effect). Two slight modifications are made to each of the component networks to better interact with ReLU. First, two neurons are used for the sum so that a negative component may be maintained in the case of $x^{3}-x$, which would otherwise be zeroed out by ReLU. Second, the triangular functions are generated upside-down, preventing ReLU from zeroing the function outside of the approximation region.
@@ 6 Concluding Remarks
This paper focused on exploiting the potential computational complexity advantages neural networks offer for the problem of efficiently learning nonlinear functions; in particular, compelling ReLU networks to approximate functions with exponential accuracy as network depth is linearly increased. Our results showed improvements of one to several orders of magnitude in using our initialization and pretraining strategy to train ReLU networks to learn various nonlinear functions, including non-convex and multi-dimensional functions. This finding is particularly powerful since random initialization and gradient descent are not likely to produce an efficient solution on their own, even if it can be proven to exist in the set of sufficiently sized ReLU networks. Although we proposed one preliminary strategy for extending our network for higher-dimensional functions, this strategy has limitations, such as leaving many weights to be filled in by gradient descent (which it generally does not do effectively). We anticipate continuing investigating strategies for improving the generalizability of our work, which we believe is critically important since it offers the possibility of an exponentially more accurate drop-in replacement of linear layers in any architecture. We are hopeful that future works by our group and others will help illuminate a complete theory for harnessing the potential exponential power of depth in ReLU or even other networks.

### v5

@@ 1 Introduction
Beyond complementary advances in areas like hardware, storage, and networking, the success of neural networks is primarily due to their ability to efficiently capture and represent nonlinear functions (Gibou et al., 2019). In a neural network, the goal of an activation function is to introduce nonlinearity between the network’s layers so that the network does not simplify to a single linear function. The rectified linear unit (ReLU) has a unique interpretation in this regard. Since it either deactivates a neuron or acts as an identity, the resulting transformation on each individual input remains linear. However, each possible configuration of active and inactive neurons can produce a unique linear transformation over a particular region of input space. The number of these activation patterns and their corresponding linear regions provides a way to measure the expressivity of a ReLU network [^1] See Appendix A.2 for definitions of terms like linear regions, activation patterns, and activation regions. and can theoretically scale exponentially with the depth of the network (Montufar et al., 2014a; Serra et al., 2018). Hence deep architectures may outperform shallow ones.
@@
The present work aims to begin eliminating these inefficiencies, starting in a simple one-dimensional setting. Drawing inspiration from existing theoretical ReLU constructions, our novel contributions include a special reparameterization of a ReLU network that forces it to maintain an exponential number of activation patterns over the input domain. We then demonstrate a novel pretraining strategy, which trains these derived parameters before manipulating the underlying matrix weights. This allows the network to discover solutions that are more accurate and unlikely to be found otherwise. Our technique directly rectifies the issues raised by Hanin & Rolnick (2019). Firstly, we initialize immediately with an exponential number of linear regions, which would not be expected to happen otherwise. Secondly, we are not reliant on gradient descent to “discover” new regions; we only need to maintain the existing ones, which is already guaranteed by the reparameterization during the pretraining stage. Additionally, this pretraining step can avoid areas of the loss landscape where unassisted gradient descent might make short-sighted optimizations that eliminate activation regions. Our results demonstrate that minimizing the reliance of network training on unassisted gradient descent can reliably produce error values orders of magnitude lower than a traditionally-trained network of equal size. Although preliminary numerical demonstrations in this theoretical study pertain to relatively low-dimensional functions, the paper concludes with our views on extending these theoretical exponential benefits for ReLU networks to arbitrary smooth functions with arbitrary dimensionality, which would have significant practical utility.
@@ 2 Related Work
This work is primarily concerned with a novel training methodology, but it also possesses a significant approximation theoretic component. In one dimension, where we initially focus, our work is a generalization of the approximation to $x^{2}$ we review in this section. The reparameterization we employ modifies this method to become trainable to represent other convex one-dimensional functions, and then converts that result back into a matrix representation. The goal of this method is not to re-express the full set of neural network weights, but to constrain them into efficient patterns.
@@
2.1 Function Approximation
Infinitely wide neural networks are known to be universal function approximators, even with only one hidden layer (Hornik et al., 1989; Cybenko, 1989). Infinitely deep networks of fixed width are universal approximators as well (Lu et al., 2017; Hanin, 2019). In finite cases, one may study trade-offs between width and depth to assess a network’s ability to approximate (learn) a function.
@@ 2.1 Function Approximation
The dilated triangular waveforms produced in this manner are not particularly useful on their own. Their oscillations quickly become excessively rapid, and their derviatives do not exist everywhere (especially in the infinite depth limit). But these problems can be rectified by taking a sum over the layers of a network. Yarotsky (2017) and Liang & Srikant (2016) construct $y=x^{2}$ on $[0,1]$ with exponential accuracy using symmetric triangle waves. To produce their approximation, one begins with $f_{0}(x)=x$, then computes $f_{1}(x)=f_{0}(x)-T(x)/4$, $f_{2}(x)=f_{1}(x)-T(T(x))/16$, $f_{3}(x)=f_{2}(x)-T(T(T(x)))/64$, and so forth, as pictured in Figure 1. As these successive approximations are computed, Figure 1 plots their convergence to $x^{2}$, as well as the convergence of the derivative to $2x$. Our reparameterization generalizes this approximation to use non-symmetric triangle waves to approximate a wider class of convex, differentiable, one-dimensional functions.
@@
The $x^{2}$ approximation is used by other theoretical works as a building block to guarantee exponential convergence rates in more complex systems. Perekrestenko et al. (2018) construct a multiplication gate via the identity $(x+y)^{2}=x^{2}+y^{2}+2xy$. The squared terms can all be moved to one side, expressing the product $xy$ as a linear combination of squared terms. They then further assemble these multiplication gates into an interpolating polynomial, which can have an exponentially decreasing error when the interpolation points are chosen to be the Chebyshev nodes. Polynomial interpolation does not scale well into high dimensions, so this and papers with similar approaches will usually come with restrictions that limit function complexity: Wang et al. (2018) requires low input dimension, Montanelli et al. (2020) uses band limiting, and Chen et al. (2019) approximates low-dimensional manifolds. These works all make use of a fixed representation of $x^{2}$. If our networks were substituted in for the $x^{2}$ approximation, these works would provide theoretical guarantees about the capabilities of the resulting model. Even though the approximation rates will not scale well with input dimension, they serve as a bound that can be improved upon. In Section 5 we further elaborate on how to use our networks to represent higher-dimensional or non-convex functions.
@@
Other works focus on showing how ReLU networks can encode and subsequently surpass traditional approximation methods (Lu et al., 2021; Daubechies et al., 2022), including spline-type methods (Eckle & Schmidt-Hieber, 2019). Interestingly, certain fundamental themes from above like composition, triangles, or squaring are still present. Another interesting comparison of the present work is to Ivanova & Kubat (1995), which uses decision trees as a means to initialize sigmoid neural networks for classification. Similar to the spirit of our work, which restricts parameterizations of ReLU networks, Elbrächter et al. (2019) explores theoretical aspects of the conditionining of ReLU network training and provides constructive results for a parameterization space that is well-conditioned. Chen & Ge (2024) present a creative approach where they explore reparameterizing the direction of weight vectors using hyperspherical coordinates to improve training dynamics. Unlike their reparameterization, ours will restrict the network’s expressivity in order to prevent it from learning inefficient weight patterns. Lastly, Park et al. (2021) approaches the problem of linear region maximization from an information theory perspective and uses a loss penalty rather than a reparameterization to increase the number of linear regions.
2.2 Neural Network Initialization
@@
Our work seeks to improve network initialization by making use of explicit theoretical constructs. This stands in sharp contrast the current standard approach, which treats neurons homogeneously. Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot & Bengio, 2010). They use weight values that are sampled from distributions defined by the input and output dimension of each layer. Aside from sub-optimal approximation power associated with random weights, a common issue is that the initial weights and biases in a ReLU network can cause every neuron in a particular layer to output a negative value. The ReLU activation then sets the output of that layer to 0, blocking any gradient updates. This is referred to as the dying ReLU phenomenon (Qi et al., 2024; Nag et al., 2023). Worryingly, as depth goes to infinity, the dying ReLU phenomenon becomes increasingly likely (Lu et al., 2020). Several papers propose solutions: Shin & Karniadakis (2020) use a data-dependent initialization, while Singh & Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed. We observed during our experiments that RAAI greatly reduces, but does not eliminate the likelihood of dying ReLU. Our approach enforces a specific network structure that does not collapse in this manner.
@@ 3 Initialization and Pretraining Construction
Naively using the output of one triangle generator (from Figure 2) as input to the next would form a $1\times 2\times 1\times 2\times 1$ shape, but this is unnecessarily deep. We can replicate the one-dimensional function composition in the hidden layers on the left side of Figure 2 by using weight sharing instead. Any outgoing weight from $t_{1}$ or $t_{2}$ is shared; every neuron taking in a triangle wave as input does so by combining $t_{1}$ and $t_{2}$ in the same proportion. In this way, we can avoid having to use extra intermediate neurons. This is evident in Figure 3 where the waves output by $t_{1}$ and $t_{2}$ neurons posses identical slopes in regions where both neurons are active.
@@ 3.1 Differentiable Model Output
Given the rapid oscillations of the triangle waves formed at each layer, the network will output a fractal with many choices of scaling parameters. This would be poorly predictive of unseen data points generated from a smooth curve. Our main mathematical result addresses this issue by forcing differentiability of the network output. We do so by making the sum of the derivatives of the triangle waves converge. For this class of functions, the peak locations of the triangle waves uniquely determine the scales with which to sum them. For a broader mathematical discussion, as well as sufficient conditions for differentiability in the limit, see the appendix.
@@
$F^{\prime}(x)=\sum_{i=0}^{\infty}s_{i}W^{\prime}_{i}(x)$ and is continuous on $[0,1]$ only if the scaling coefficients $s_{i}$ are selected based on the triangle peaks $a_{i}$ according to:
@@ 4 One-Dimensional Experiments
The goal of deep learning is to train a network to approximate a (generally unknown) nonlinear function. Accordingly, in this section, we implement our method to learn several convex one-dimensional curves. These nonlinear functions are known—so that we can measure error—and we remark and demonstrate results on more difficult functions in Section 5, Section 6, and Appendix A.10. The aim of these experiments is twofold: (1) we would like to determine how to learn the most effective function representations possible, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data. To demonstrate that our networks can learn function representations that better utilize depth, we benchmark against PyTorch’s (Paszke et al., 2019) default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh & Sreejith (2021), and produce errors that are orders of magnitude lower than both. For our experimental networks, we use a common set of initialization points where the triangle peaks and scaling parameters are chosen according to our main theorem (Equation 4). We compare the effect of reparameterized pretraining against skipping immediately to standard gradient descent (from a reparameterized initialization as in Section 3). We also compare training the scaling parameters freely during pretraining, instead of choosing them to achieve differentiability. The intent of both comparisons is to see if pretraining and further differentiability constraints facilitate smoother navigation of the loss landscape. Lastly, we conduct a second round of tests to determine if pretrained networks display an enhanced predictive capacity on unseen data points, as might be expected if they can leverage greater nonlinearity in their outputs.
@@ 4.2 Numerical Results
Default Network (Kaiming) $2.11\times 10^{-5}$ $2.19\times 10^{-5}$ $7.20\times 10^{-2}$ $2.82\times 10^{-2}$
@@
Default Network (Kaiming) $4.50\times 10^{-5}$ $5.75\times 10^{-5}$ $1.15\times 10^{-1}$ $1.96\times 10^{-1}$
@@
Default Network (Kaiming) $2.41\times 10^{-4}$ $2.14\times 10^{-3}$ $2.27\times 10^{-5}$ $1.60\times 10^{-4}$
@@
Our first set of results are shown in Tables 1 and 2, wherein we observe several important trends. First, the worst performing networks are the Default Networks that rely on randomized (Kaiming) initialization. Even the networks that forgo pretraining benefit from initializing with many activation regions. When pretraining constraints are used, they are able to steer gradient descent to the best solutions, resulting in reductions in minimum error of three orders of magnitude over default networks. Pretraining with differentiability enforced also closes the gap between the minimum and mean errors compared to other setups. This indicates that these loss landscapes are indeed the most reliable to traverse. Enforcing differentiability during pretraining can impart a bias towards smoother solutions during subsequent unassisted gradient descent.
@@
Figure 4: Comparison between standard Kaiming initialization/gradient descent and pretraining with differentiablity enforced. Using more linear regions allows the curve to better predict the test points.
@@ 5 Extension to Non-Convex Functions and Higher Dimensions
The same approach can address non-convex functions since many non-convex functions (including every continuous function with a bounded second derivative (Hessian)) can be expressed as the difference of two convex functions (see Zhang et al. (2018), which also describes how every ReLU network can be decomposed into a difference of two piecewise linear convex functions). We do not claim that two copies of our network are sufficient for learning every pathological function, but preliminary results in this section are surprisingly promising. In fact, we highly encourage reviewing Appendix A.10, which demonstrates our method on real-world datasets and problems with up to eight input dimensions.
@@
Figures 7 and 8 show learning the non-convex function $y=x^{3}-x$ and the two-dimensional function $z=r^{3}$ on $[-1,1]$, each by using two of our convex function blocks. Rather than directly applying our networks as activation functions, we instead encode each of their layers as a block diagonal structure in one combined matrix (which has the same effect). Two slight modifications are made to each of the component networks to better interact with ReLU. First, two neurons are used for the sum so that a negative component may be maintained in the case of $x^{3}-x$, which would otherwise be zeroed out by ReLU. Second, the triangular functions are generated upside-down, preventing ReLU from zeroing the function outside of the approximation region.
@@ 6 Concluding Remarks
This paper focused on exploiting the potential computational complexity advantages neural networks offer for the problem of efficiently learning nonlinear functions; in particular, compelling ReLU networks to approximate functions with exponential accuracy as network depth is linearly increased. Our results showed improvements of one to several orders of magnitude in using our initialization and pretraining strategy to train ReLU networks to learn various nonlinear functions, including non-convex and multi-dimensional functions. This finding is particularly powerful since random initialization and gradient descent are not likely to produce an efficient solution on their own, even if it can be proven to exist in the set of sufficiently sized ReLU networks. Although we proposed one preliminary strategy for extending our network for higher-dimensional functions, this strategy has limitations, such as leaving many weights to be filled in by gradient descent (which it generally does not do effectively). We anticipate continuing investigating strategies for improving the generalizability of our work, which we believe is critically important since it offers the possibility of an exponentially more accurate drop-in replacement of linear layers in any architecture. We note that there are other metrics for evaluating a neural network’s expressivity (e.g., Raghu et al. (2017)), which would be interesting to measure for our networks in future work. Overall, we are hopeful that future works by our group and others will help illuminate a complete theory for harnessing the potential exponential power of depth in ReLU and other classes of neural networks.